# MuRating: A High Quality Data Selecting Approach to Multilingual Large Language Model Pretraining

### v1

@@ Abstract
Data quality is a critical driver of large language model performance, yet existing model-based selection methods focus almost exclusively on English. We introduce MuRating, a scalable framework that transfers high-quality English data-quality signals into a single rater for 17 target languages. MuRating aggregates multiple English “raters” via pairwise comparisons to learn unified document-quality scores, then projects these judgments through translation to train a multilingual evaluator on monolingual, cross-lingual, and parallel text pairs. Applied to web data, MuRating selects balanced subsets of English and multilingual content to pretrain a 1.2 B-parameter LLaMA model. Compared to strong baselines, including QuRater, AskLLM, DCLM and so on, our approach boosts average accuracy on both English benchmarks and multilingual evaluations, with especially large gains on knowledge-intensive tasks. We further analyze translation fidelity, selection biases, and underrepresentation of narrative material, outlining directions for future work.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of tasks, and recent studies have consistently emphasized the critical role of high-quality pretraining data in driving these advances brown2020language ; rae2021scaling ; chowdhery2023palm . To improve data quality, various strategies have been adopted, such as deduplication lee2021deduplicating ; abbas2023semdedup , heuristic and rule-based filtering rae2021scaling ; penedo2024fineweb , and domain-aware sampling xie2023doremi ; shen2023slimpajama . While effective, these methods often rely heavily on manual heuristics and domain expertise, lacking a unified or principled framework for evaluating and selecting pretraining data. Moreover, they are typically applied as pre-defined or post-hoc filters, limiting their adaptability to downstream performance. In response, model-based data selection approaches have emerged, aiming to learn data quality judgments from examples or auxiliary supervision. These methods utilize different model architectures and data selection criteria. For instance, DCLM li2024datacomp trains a FastText classifier joulin2016bag using high-quality samples from OH2.5 and Reddit ELI5 as positive supervision, while treating Common Crawl web data as negatives. Other approaches such as AskLLM, QuRater, and the FineWeb-Edu classifier sachdeva2024train ; wettig2024qurating ; lozhkov2024fineweb-edu employ prompt-based evaluation criteria using various LLMs to assess the quality of input samples.
@@
Data selection beyond English remains a profound challenge fang-etal-2017-learning ; laurenccon2022bigscience . While model-based data selection methods have demonstrated effectiveness in improving training quality, they have been developed almost entirely for English, leaving a critical gap in multilingual data quality assessment. As LLMs are increasingly applied in diverse linguistic contexts, there is a growing need for selection strategies that extend beyond English. A recent attempt Fineweb2-hq messmer2025enhancing train language-specific raters using benchmark datasets as positive supervision and general pretraining corpora as negatives, following a strategy similar to DCLM. However, this approach trains separate raters per language, demanding large, human-curated positive examples that simply do not exist for many low-resource languages. Additionally, using benchmark-derived data poses a risk of contamination.
@@
In this work, we introduce MuRating, a two-stage, translation-and-pairwise framework for multilingual data-quality estimation. We begin by aggregating multiple state-of-the-art English raters via majority-vote pairwise comparisons, fitting a Bradley–Terry model bradley1952rank to learn a single, unified quality scorer. Next, we translate scored English document pairs into each of 17 target languages and construct monolingual, cross-lingual, and parallel pairs—projecting original preference labels onto translated comparisons and assigning neutral labels to parallel translations. This design yields one multilingual evaluator that preserves English-derived quality signals while remaining language-agnostic.
@@
We apply MuRating framework to fine-tune a MuRater model to annotate 1.5 trillion English and 3 trillion multilingual web tokens, selecting 200 billion English and 300 billion multilingual tokens for pretraining a 1.2 B-parameter LLaMA model. Compared to strong baselines—uniform sampling (+50 % data), QuRater, AskLLM, DCLM—our selection yields an average gain of 1 to 3.4 points on twelve English benchmarks (including ARC, HellaSwag, MMLU) and 1.8 points on a diverse multilingual suite (XCOPA, XNLI, Flores, plus native MMLU-variants). We further assess translation fidelity via human evaluation, examine the impact of cross-lingual and parallel data, and compare different score transfer approaches.
Our contributions are as follows:
@@
  • Scalable pretraining gains. MuRating’s selected 500 billion-token corpus yields substantial improvements over state-of-the-art baselines on both English and multilingual LLM benchmarks.
  • Analysis and open resources. We assess translation fidelity, address selection biases, and will release our prompts, code, and data to facilitate further research.
@@ 2 Related Work
Data Selection. Data selection is essential in constructing high-quality pretraining corpora for LLMs and typically falls into three main categories: deduplication, heuristic-based filtering, and LLM-guided quality evaluation. Deduplication, applied early, removes exact or near-duplicate documents to reduce redundancy and improve generalization lee2021deduplicating . More advanced fuzzy and semantic methods filter syntactically or semantically similar content jiang2022fuzzydedup ; abbas2023semdedup , which is crucial at scale to avoid training instability and performance degradation xue2023repeat ; rae2021scaling .
@@
Heuristic filtering uses rules or lightweight models to exclude low-quality text, such as short, repetitive, or toxic content laurenccon2022bigscience ; weber2024redpajama ; penedo2023refinedweb ; soldaini2024dolma . While handcrafted heuristics can be effective, they often have limited generalization and inefficiency, prompting the use of simple classifiers, perplexity scores, or importance sampling chowdhery2023palm ; touvron2023llama ; xie2023data ; muennighoff2023scaling . However, these approaches may unintentionally favor simplistic or repetitive content, which can diminish the diversity and informativeness of the dataset.
@@
In contrast, LLM-guided quality scoring directly leverages language models to evaluate data along dimensions like factuality and coherence gunasekar2023textbooks ; sachdeva2024train . Frameworks such as QuRating and FineWeb-Edu prioritize educational content using multi-criteria assessment wettig2024qurating ; lozhkov2024fineweb-edu , while Dataman peng2025dataman and FIRE xu2025fire extend this to domain-aware or reliability-sensitive filtering. Despite their advancements, recent approaches depend heavily on GPT-style judgments, potentially introducing model-specific biases.
@@
Multilingual Pretraining. Efforts to construct multilingual datasets for multilingual LLM pretraining have followed similar strategies to those used for English, incorporating deduplication and heuristic-based filtering techniques. Prominent corpora such as mC4 xue2020mt5 , RedPajama weber2024redpajama , CulturalX nguyen2023culturax , HPTL de2024new , and FineWeb-2 penedo2024fineweb leverage these methods to scale multilingual resources, ranging from a few dozen to thousands of languages, significantly enhancing cross-lingual performance in multilingual LLM pretraining. To mitigate the issue of data scarcity for low-resource languages, TransWeb-Edu wang2024multilingual addresses this by translating high-quality English data into multiple languages. However, there has been limited research on model-based data selection in the context of multilingual LLM pretraining. Recently, an initial approach messmer2025enhancing introduced a model-based selection method to refine the FineWeb-2 dataset by training language-specific classifiers, using multilingual benchmark datasets as positive examples and web corpus data as negative examples. However, this method relies on the availability of high-quality samples from existing multilingual benchmarks, which presents scalability challenges and limits its applicability to low-resource or underrepresented languages. Additionally, this approach may risk contaminating downstream evaluation tasks with biased data.
@@ 3 Methodology
Our approach consists of two stages: (1) consolidating multiple English-language quality raters into a single, unified scorer via pairwise comparisons, and (2) transferring that scorer to a multilingual setting through translation-based alignment and cross-lingual regularization. We introduce a two-step method: first, we integrate existing English corpus quality raters; second, we transfer their rating capability to a multilingual setting.
3.1 Integration of English Raters
@@
We define a binary preference for each rater: if $S_{A}^{n}>S_{B}^{n}$, we consider that rater $n$ prefers text $t_{A}$ over $t_{B}$, and vice versa. Based on the collective preferences of all raters, we compute an empirical confidence score $P_{A>B}$ indicating how likely $t_{A}$ is preferred over $t_{B}$:
$P_{A>B}=\frac{1}{|N|}\sum_{n\in N}\mathbb{I}[S_{A}^{n}>S_{B}^{n}],\quad P_{A>B}\in[0,1],$ (1)
where $\mathbb{I}[\cdot]$ is the indicator function that equals 1 when the condition is true and 0 otherwise. This score quantifies the relative preference strength of $t_{A}$ over $t_{B}$ across all raters.
@@
where, $\sigma(\cdot)$ denotes the sigmoid function, and $p_{B\succ A}=1-P_{A>B}$ is the empirical probability that $t_{B}$ is preferred over $t_{A}$. This formulation encourages the model to assign higher scores to texts that are consistently preferred in the pairwise judgments.
After training, the model outputs a scalar quality rating for each document. These scores are treated as logits over the dataset and are used for quality-based sampling, where a subset of high-quality texts is selected based on their relative scores.
@@ 3.2.1 Translation-Based Alignment of Multilingual Preferences
This assumption is based on the premise that translation preserves both the semantic content and the relative quality between text pairs. Prior work QuRating wettig2024qurating highlights that pairwise comparisons offer increased stability when evaluating text quality. In multilingual settings, pointwise scoring—where absolute quality scores are assigned to individual texts—tends to be more susceptible to subtle changes in tone or phrasing introduced during translation, which can compromise the consistency of the supervision signal. In contrast, pairwise supervision is inherently more robust to such translation-induced variations. As long as the relative ranking between the texts remains consistent (i.e., $t_{A}^{en}$ continues to be preferred over $t_{B}^{en}$ after translation), the corresponding translated pair $(t_{A}^{m},t_{B}^{m})$ remains a valid training example. This robustness makes pairwise comparisons a more reliable and effective framework for training quality evaluation models in multilingual contexts.
@@ 3.3 Cross-Lingual and Language-Agnostic Alignment
While the previous section addressed only in-language supervision—i.e., training on text pairs $(t_{A}^{m},t_{B}^{m})$ where both documents are in the same language $m$—this setup alone is insufficient to guarantee language-agnostic scoring behavior. To promote consistency in quality assessments across languages, we augment our training dataset with both cross-lingual and parallel text pairs.
For cross-lingual supervision, we generate mixed-language pairs by translating $t_{A}$ and $t_{B}$ into different target languages, resulting in pairs of the form $(t_{A}^{m},t_{B}^{m^{\prime}})$ with $m\neq m^{\prime}$. The original English pairwise preference score is then transferred to these cross-lingual pairs by assuming $P_{A^{m}>B^{m^{\prime}}}\approx P_{A^{en}>B^{en}}$.
In addition, we introduce parallel pairs to explicitly regularize the model’s behavior on semantically equivalent content across languages. Given a text $t_{A}^{m}$ and its direct translation $t_{A}^{m^{\prime}}$ into another language $m^{\prime}$, we form the pair $(t_{A}^{m},t_{A}^{m^{\prime}})$ and assign a neutral preference score, i.e., $P_{A^{m}>A^{m^{\prime}}}\approx 0.5$. This reflects the expectation that both texts, despite being in different languages, convey identical semantic meaning and should be treated as equally quality.
@@
$\mathcal{L}_{\text{parallel}}=\underset{(t_{A},t_{A^{{}^{\prime}}})\in\mathcal{J^{{}^{\prime}}}}{\mathbb{E}}\Bigg{[}-\log\sigma\left(s_{\theta}(t_{A}^{m})-s_{\theta}(t_{A}^{m^{\prime}})\right)-\log\sigma\left(s_{\theta}(t_{A}^{m^{\prime}})-s_{\theta}(t_{A}^{m})\right)\Bigg{]},$ (3)
@@ 3.3.1 Multilingual Rater Objective
where $\lambda$ is a tunable hyperparameter balancing cross-lingual consistency and discrimination. This joint training approach allows us to construct a multilingual quality rater that is robust, consistent across languages, and sensitive to relative quality differences.
@@ 3.3.2 Training the Rater Model
To build a high-quality multilingual rater, we sample 300,000 English text pairs and annotate them using four rating methods. For GPT-4o-based annotation, we prompt the model in both directions—$(t_{A},t_{B})$ and $(t_{B},t_{A})$—multiple times to mitigate order bias. The final confidence score $P_{A>B}$ is computed by averaging predicted preference probabilities. For other raters (AskLLM, FineWeb-Edu-Classifier, DCLM), we collect individual scores and then integrate pairwise preferences using Equation 1.
We translate English pairs into 150,000 monolingual and 150,000 cross-lingual pairs, where languages portion is balanced across languages. Additionally, 75,000 texts are translated into two different languages to form parallel pairs $(t_{A}^{m},t_{A}^{m^{\prime}})$, which are assigned a neutral score of $P_{A>A^{\prime}}=0.5$ to promote consistency across languages for semantically equivalent content.
The final MuRater training set includes 75,000 English, 150,000 monolingual, 150,000 cross-lingual, and 75,000 parallel pairs. We adopt QuRater’s training setup wettig2024qurating , applying a confidence margin to all but the parallel examples.
We fine-tune an encoder-based model following the BGE-M3 architecture bge-m3 , adding a linear head to predict quality ratings. The resulting rater achieves over 93% accuracy on held-out and 97% on held-in pairwise judgments, demonstrating strong multilingual preference modeling. Implementation details, including tokenizer settings and hyperparameters, are provided in Appendix A.

### v2

@@ Abstract
Data quality is a critical driver of large language model performance, yet existing model-based selection methods focus almost exclusively on English, neglecting other languages that are essential in the training mix for multilingual LLMs. We introduce MuRating, a scalable framework that transfers high‐quality English data‐quality signals into a multilingual autorater, capable of handling 17 languages. MuRating aggregates multiple English autoraters via pairwise comparisons to learn unified document quality scores, then projects these judgments through translation to train a multilingual evaluator on monolingual, cross‐lingual, and parallel text pairs. Applied to web data, MuRating selects balanced subsets of English and multilingual content to pretrain LLaMA-architecture models of 1.2B and 7B parameters. Compared to strong baselines, including QuRater, FineWeb2-HQ, AskLLM, DCLM, our approach increases average accuracy on both English benchmarks and multilingual evaluations. Extensive analyses further validate that pairwise training provides greater stability and robustness than pointwise scoring, underscoring the effectiveness of MuRating as a general multilingual data-selection framework.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of tasks, and recent studies have consistently emphasized the critical role of high-quality pretraining data in driving these advances brown2020language; rae2021scaling; chowdhery2023palm. To improve data quality, various strategies have been adopted, such as deduplication lee2021deduplicating; abbas2023semdedup, heuristic and rule-based filtering rae2021scaling; penedo2024fineweb, and domain-aware sampling xie2023doremi; shen2023slimpajama. While effective, these methods often rely heavily on manual heuristics and domain expertise, lacking a unified or principled framework for evaluating and selecting pretraining data. Moreover, they are typically applied as pre-defined or post-hoc filters, limiting their adaptability to downstream performance. In response, model-based data selection approaches have emerged, aiming to learn data quality judgments from examples or auxiliary supervision. These methods utilize different model architectures and data selection criteria. For instance, DCLM li2024datacomp trains a FastText classifier joulin2016bag using high-quality samples from OH2.5 and Reddit ELI5 as positive supervision, while treating Common Crawl web data as negatives. Other approaches such as AskLLM sachdeva2024train, QuRater wettig2024qurating, and the FineWeb-Edu lozhkov2024fineweb-edu classifier employ prompt-based evaluation criteria using various LLMs to assess the quality of input samples.
@@
Data selection beyond English remains an important challenge fang-etal-2017-learning; laurenccon2022bigscience. While model-based data selection methods have demonstrated effectiveness in improving training quality, they have been developed almost entirely for English and are not explicitly designed or validated for non-English languages, leaving a critical gap in multilingual data quality assessment. As LLMs are increasingly applied in diverse linguistic contexts, there is a growing need for selection strategies that extend beyond English. A recent attempt Fineweb2-HQ messmer2025enhancing train language-specific raters using benchmark datasets as positive supervision and general pretraining corpora as negatives, following a strategy similar to DCLM li2024datacomp. However, this approach uses benchmark-derived data, posing a risk of test set contamination.
@@
In this work, we introduce MuRating, a two-stage, translation-and-pairwise framework for multilingual data-quality estimation. MuRating begins by aggregating multiple state-of-the-art English raters via majority-vote pairwise comparisons, fitting a Bradley–Terry model bradley1952rank to learn a single, unified quality scorer. Next, it translates scored English document pairs into each of 17 target languages and construct monolingual, cross-lingual, and parallel pairs—projecting original preference labels onto translated comparisons and assigning neutral labels to parallel translations. Here, parallel pairs consist of identical content translated into two different languages, while cross-lingual pairs involve distinct texts written in different languages. This design yields one multilingual evaluator that preserves English-derived quality signals while remaining language-agnostic.
@@
We apply MuRating framework to fine-tune a MuRater model to annotate English and multilingual web documents and select top 10% data to pretrain LLaMA-architecture grattafiori2024llama models of 1.2B and 7B parameters for validation. Compared to strong baselines—uniform sampling with 50% more data, QuRater wettig2024qurating, AskLLM sachdeva2024train, FineWeb2-HQ messmer2025enhancing—our selection yields an average gain of 1 to 3.4 points on twelve English benchmarks and 1.8 points on a diverse multilingual suite. We further assess translation fidelity via human evaluation, examine the impact of cross-lingual and parallel data, and compare different score transfer approaches.
Our contributions are as follows: [^1] Our code is available at: https://github.com/aialt/MuRater.
@@
  • Scalable pretraining gains. The results from both 1.2B and 7B model experiments demonstrate significant gains over state-of-the-art baselines across English and multilingual LLM benchmarks.
@@ 2 Related Work
Data Selection. Data selection is essential in constructing high-quality pretraining corpora for LLMs and typically falls into three main categories: deduplication, heuristic-based filtering, and LLM-guided quality evaluation. Early-stage deduplication removes exact or near-duplicate documents to minimize redundancy and enhance model generalization lee2021deduplicating. More advanced fuzzy and semantic methods filter syntactically or semantically similar content jiang2022fuzzydedup; abbas2023semdedup, which is crucial at scale to avoid training instability and performance degradation xue2023repeat; rae2021scaling.
@@
Heuristic filtering uses rules or lightweight models to exclude low-quality text, such as short, repetitive, or toxic content laurenccon2022bigscience; weber2024redpajama; penedo2023refinedweb; soldaini2024dolma. While handcrafted heuristics can be effective, they often have limited generalization and inefficiency, prompting the use of simple classifiers, perplexity scores, or importance sampling chowdhery2023palm; touvron2023llama; xie2023data; muennighoff2023scaling; li-etal-2024-quantity; li-etal-2024-superfiltering. However, these approaches may unintentionally favor simplistic or repetitive content, which can diminish the diversity and informativeness of the dataset.
@@
In contrast, LLM-guided quality scoring directly leverages language models to evaluate data along dimensions like factuality and coherence gunasekar2023textbooks; sachdeva2024train; li-etal-2024-selective. Frameworks such as QuRating wettig2024qurating and FineWeb-Edu lozhkov2024fineweb-edu prioritize educational content using multi-criteria assessment , while Dataman peng2025dataman and FIRE xu2025fire extend this to domain-aware or reliability-sensitive filtering. Despite their advancement, recent approaches depend heavily on GPT-style judgments, potentially introducing model-specific biases.
@@
Multilingual Pretraining. Efforts to construct multilingual datasets for multilingual LLM pretraining have followed similar strategies to those used for English, incorporating deduplication and heuristic-based filtering techniques. Prominent corpora such as mC4 xue2020mt5, RedPajama weber2024redpajama, CulturalX nguyen2023culturax, HPTL de2024new, and FineWeb-2 penedo2025fineweb2pipelinescale leverage these methods to scale multilingual resources, ranging from a few dozen to thousands of languages, significantly enhancing cross-lingual performance in multilingual LLM pretraining. To mitigate the issue of data scarcity for low-resource languages, TransWeb-Edu wang2024multilingual addresses this by translating high-quality English data into multiple languages. In addition, EMMA-X guo2023emma introduces an EM-inspired framework that jointly learns cross-lingual semantic alignment and sentence representations from large-scale non-parallel multilingual data, while CLIMB guo2025exploring enhances multilingual capability by dynamically adjusting the data-mix ratio across languages. However, research on model-based data selection for multilingual LLM pretraining remains limited. Recently, an initial approach messmer2025enhancing introduced a model-based selection method to refine the FineWeb-2 dataset by training language-specific classifiers, using multilingual benchmark datasets as positive examples and web corpus data as negative examples. However, this method relies on the availability of high-quality samples from existing multilingual benchmarks, which may risk contaminating downstream evaluation tasks with biased data.
@@ 3 Methodology
Our approach consists of two stages: (1) consolidating multiple English‐language quality raters into a single, unified scorer via pairwise comparisons, and (2) transferring the scorer to a multilingual setting through translation‐based alignment and cross-lingual regularization. We introduce a two-step method: first, we integrate existing English corpus quality raters; second, we transfer their rating capability to a multilingual setting.
3.1 Integration of English AutoRaters
@@
We define a binary preference for each rater: if $S_{A}^{n}>S_{B}^{n}$, we consider that rater $n$ prefers text $t_{A}$ over $t_{B}$, and vice versa. If the two scores are nearly identical (i.e., $|S_{A}^{n}-S_{B}^{n}|<\epsilon$), we treat the preference as ambiguous and discard the pair from the training dataset. Based on the remaining valid preferences from all raters, we compute an empirical confidence score $P_{A>B}$ indicating how likely $t_{A}$ is preferred over $t_{B}$:
$P_{A>B}=\frac{1}{|N|}\sum_{n\in N}\mathbb{I}[S_{A}^{n}>S_{B}^{n}],\quad P_{A>B}\in[0,1],\quad\text{where }|S_{A}^{n}-S_{B}^{n}|\geq\epsilon.$ (1)
where $\mathbb{I}[\cdot]$ is the indicator function that equals 1 when the condition is true and 0 otherwise. if This score quantifies the relative preference strength of $t_{A}$ over $t_{B}$ across all raters.
@@
where $\sigma(\cdot)$ denotes the sigmoid function, and $p_{B\succ A}=1-P_{A>B}$ is the empirical probability that $t_{B}$ is preferred over $t_{A}$. This formulation encourages the model to assign higher scores to texts that are consistently preferred in the pairwise judgments.
After training, the model outputs a single scalar score representing the quality of each document. These scores are treated as logits over the dataset and are used for quality-based sampling, where a subset of high-quality texts is selected based on their relative scores.
@@ 3.2.1 Translation-Based Alignment of Multilingual Preferences
We also experimented with the reverse approach—translating multilingual pairs into English, scoring them in English, and then projecting the scores back to the corresponding multilingual pairs. A comparison between the two setups is presented in the experiment section 4.2.1.
This assumption is based on the premise that translation preserves both the semantic content and the relative quality between text pairs. Prior work QuRating wettig2024qurating highlights that pairwise comparisons offer increased stability when evaluating text quality. In multilingual settings, pointwise scoring—where absolute quality scores are assigned to individual texts—is more susceptible to subtle changes in tone or phrasing introduced during translation, which can compromise the consistency of the supervision signal. In contrast, pairwise supervision is inherently more robust to such translation-induced variations. As long as the relative ranking between the texts remains consistent (i.e., $t_{A}^{en}$ continues to be preferred over $t_{B}^{en}$ after translation), the corresponding translated pair $(t_{A}^{m},t_{B}^{m})$ remains a valid training example. This robustness makes pairwise comparisons a more reliable and effective framework for training quality evaluation models in multilingual contexts.
@@ 3.3 Cross-Lingual and Language-Agnostic Alignment
While the previous section addressed only in-language supervision—i.e., training on text pairs $(t_{A}^{m},t_{B}^{m})$ where both documents are in the same language mm—this setup alone is insufficient to guarantee language-agnostic scoring behavior. To promote consistency in quality assessments across languages, we augment our training dataset with both cross-lingual and parallel text pairs.
For cross-lingual pair construction, we generate mixed-language pairs by randomly translating $t_{A}$ and $t_{B}$ into different target languages, resulting in pairs of the form $(t_{A}^{m},t_{B}^{m^{\prime}})$ with $m\neq m^{\prime}$. The original English pairwise preference score is then transferred to these cross-lingual pairs by assuming $P_{A^{m}>B^{m^{\prime}}}\approx P_{A^{en}>B^{en}}$.
For parallel pair construction, we build semantically equivalent pairs to explicitly regularize the model’s behavior. Given a text $t_{A}^{m}$ and its direct translation $t_{A}^{m^{\prime}}$ into another language $m^{\prime}$, we form the pair $(t_{A}^{m},t_{A}^{m^{\prime}})$ and assign a neutral preference score, i.e., $P_{A^{m}>A^{m^{\prime}}}\approx 0.5$. This reflects the expectation that both texts, despite being in different languages, convey identical semantic meaning and should be treated as equally quality.
@@
$\mathcal{L}_{\text{parallel}}=\mathbb{E}_{(t_{A}^{m},t_{A}^{m^{\prime}})\in\mathcal{J}^{\prime}}\left[-\log\sigma\!\left(s_{\theta}(t_{A}^{m})-s_{\theta}(t_{A}^{m^{\prime}})\right)-\log\sigma\!\left(s_{\theta}(t_{A}^{m^{\prime}})-s_{\theta}(t_{A}^{m})\right)\right],$ (3)
@@ 3.3.1 Multilingual Rater Objective
where $\mathcal{L}_{\text{pairwise}}$ means the loss calcuated by equation 2 for both monolingual and cross-lingual pairs, $\lambda$ is a tunable hyperparameter balancing cross-lingual consistency and discrimination. This joint training approach allows us to construct a multilingual quality rater that is robust, consistent across languages, and sensitive to relative quality differences.
@@ 3.3.2 Training the Rater Model
[…]
We then extend these English pairs into multilingual settings using GPT-4o for translation: 150,000 monolingual pairs, 150,000 cross-lingual pairs, and 75,000 parallel pairs $(t_{A}^{m},t_{A}^{m^{\prime}})$, with language proportions balanced across all target languages. The combined dataset—comprising English, monolingual, cross-lingual, and parallel examples—forms the final MuRater training set. We adopt QuRater’s training setup wettig2024qurating, applying a confidence margin to all but the parallel examples.
We fine-tune an encoder-based model following the BGE-M3 architecture bge-m3, adding a linear head to predict quality ratings. We choose BGE-M3 bge-m3 for its strong multilingual representation ability and lightweight design, which make it well-suited for large-scale multilingual scoring. The resulting rater achieves over 93% accuracy on the validation set and 97% on the training set, demonstrating strong multilingual preference modeling. The effect of translation quality and implementation details, including tokenizer settings and hyperparameters, are provided in Appendix 6.